Determine whether the two lines are parallel, perpendicular, or neither.
3y + 21 = 2x
2y = 16 - 3x

Answer by mananth(16946)

3y + 21 = 2x
3y=2x-21
/3
y=(2/3)x-7
slope m1= (2/3)
-----
2y = 16 - 3x
/2
y=(-3/2)x +8
slope m2= (-3/2)
The slopes are negative reciprocals
m1*m2=-1
(-3/2)*(2/3)=-1
So the lines are perpendicular